# Fact-Checking the Output of Large Language Models via Token-Level Uncertainty Quantification

### v1

@@ Abstract
Large language models (LLMs) are notorious for hallucinating, i.e., producing erroneous claims in their output. Such hallucinations can be dangerous, as occasional factual inaccuracies in the generated text might be obscured by the rest of the output being generally factual, making it extremely hard for the users to spot them. Current services that leverage LLMs usually do not provide any means for detecting unreliable generations. Here, we aim to bridge this gap. In particular, we propose a novel fact-checking and hallucination detection pipeline based on token-level uncertainty quantification. Uncertainty scores leverage information encapsulated in the output of a neural network or its layers to detect unreliable predictions, and we show that they can be used to fact-check the atomic claims in the LLM output. Moreover, we present a novel token-level uncertainty quantification method that removes the impact of uncertainty about what claim to generate on the current step and what surface form to use. Our method Claim Conditioned Probability (CCP) measures only the uncertainty of particular claim value expressed by the model. Experiments on the task of biography generation demonstrate strong improvements for CCP compared to the baselines for six different LLMs and three languages. Human evaluation reveals that the fact-checking pipeline based on uncertainty quantification is competitive with a fact-checking tool that leverages external knowledge.
@@ 1 Introduction
Large language models (LLMs) have become a ubiquitous and versatile tool for addressing a variety of natural language processing (NLP) tasks. People use these models for tasks including information search Sun et al. (2023b), to ask medical questions Thirunavukarasu et al. (2023), or to generate new content Sun et al. (2023a). There has been notable shift in user behavior, indicating an increasing reliance on and trust in LLMs as primary information sources, often over traditional channels. However, a significant challenge that arises with the spread of these models is their tendency to produce “hallucinations”, i.e., factually incorrect generations featuring misleading information Bang et al. (2023); Dale et al. (2023). This is a side-effect of the way modern LLMs are designed and trained Kalai and Vempala (2023).
LLM hallucinations are a major concern because the deceptive content at the the surface-level can be very coherence and persuasive. Common examples include the creation of fictitious biographies or the assertion of unfounded claims. The danger is that a few occasional false claims might be easily obscured by a large number of factual statements, making it extremely hard for people to spot them. As hallucinations in LLM outputs are hard to eliminate completely, users of such systems could be informed via highlighting some potential caveats in the text, and this is where our approach can help.
Fact-checking is a research direction that addresses this problem. It is usually approached using complex systems that leverage external knowledge sources Guo et al. (2022); Nakov et al. (2021); Wadden et al. (2020). This introduces problems related to the incomplete nature of such sources and notable overhead in terms of storing the knowledge. We argue that information about whether a generation is a hallucination is encapsulated in the model output itself, and can be extracted using uncertainty quantification (UQ) techniques Gal et al. (2016). This avoids implementing complex and expensive fact-checking systems that require additional computational overhead and rely on external resources.
Prior work has mainly focused on quantification of uncertainty for the whole generated text and been mostly limited to tasks such as machine translation Malinin and Gales (2020), question answering Kuhn et al. (2022), and text summarization van der Poel et al. (2022). However, the need for UQ of only a portion of the generations substantially complicates the problem. We approach it by leveraging token-level uncertainty scores and aggregating them into claim-level scores. Moreover, we introduce a new token-level uncertainty score, namely claim-conditioned probability (CCP), which demonstrates confident improvements over several baselines for six LLMs and three languages.
@@
  • We propose a novel method for token-level uncertainty quantification that outperforms baselines and can be used as a plug-in in a fact-checking pipeline.
@@
  • We provide an empirical and ablation analysis of method for fact-checking of LLM generations, and find that the uncertainty scores we produce can help to spot claims with factual errors for six LLMs over three languages: English, Chinese and Arabic.
  • All code and data will be made available upon acceptance.
@@
Figure 1: Visual comparison of our proposed Claim-Conditioned Probability method to the Maximum Probability baseline. CCP accurately identifies the incorrectly specified number of awards (in red), whereas Maximum Probability erroneously highlights the claim that is actually correct.
@@ 2.1 Fact-Checking LLM Generations and Detecting Hallucinations
The problem of hallucinations has made fact-checking of LLM outputs a prominent topic in the research community, and resulted in a surge of work. Chern et al. (2023) present Factool – a task and domain agnostic framework for hallucination detection that leverages GPT for claim extraction and verification. Manakul et al. (2023) suggest to sample multiple outputs from black-box LLMs and evaluate how similar the sampled responses are using the external model. Varshney et al. (2023) detect LLM hallucinations by extracting key parts of output using an external model and estimating their uncertainty based on logits. The most uncertain parts are verified using an external knowledge source. Pan et al. (2023) propose to fact-check complex statements by decomposing them into simpler subtasks and generating reasoning programs to verify these statements. Min et al. (2023) present a methodology for evaluating long LLM-generated texts by decomposing them into simple atomic statements and further verifying them against some knowledge source.
@@
In contrast to previous work that leverages external knowledge sources for fact-checking (a database or another LLM), our work is the first to investigate token-level UQ methods for this task while having access to the LLM and its outputs only.
2.2 Uncertainty Quantification of LLM Generation
@@
UQ techniques for LLM generation can be classified into five major categories Fadeeva et al. (2023). Information-based methods leverage the probability distribution of generated tokens and usually do not require additional models. In this category, we can include methods such as perplexity Fomicheva et al. (2020), mean token entropy Fomicheva et al. (2020), point-wise mutual information Takayama and Arase (2019), and conditional point-wise mutual information van der Poel et al. (2022).
@@
Another category is based on density estimation of latent model representations. It includes Mahalanobis distance Lee et al. (2018) with various modifications Ren et al. (2023). The disadvantage of such approaches is the need for access to the LLM training data in order to fit external density models, which is infeasible for most LLMs.
@@
Ensembling and Monte Carlo dropout methods are based on the lexical diversity of multiple LLM outputs for a single query Malinin and Gales (2020); Fomicheva et al. (2020). Their main drawback is that they require many predictions, which makes them too computationally and memory intensive for practical purposes.
@@
It has recently been shown that LLMs can reflexively estimate the confidence of their generations simply by asking themselves about the truthfulness of their output Kadavath et al. (2022). This can work better than analyzing the probability distribution for the original prediction, but requires a second pass of inference, feeding the original output as a part of the query.
@@
Finally, there is a group of methods that leverages diversity of meanings that the LLM generates for a given query. This group includes semantic entropy Kuhn et al. (2022) and various scores based on the analysis of the similarity matrix between outputs Lin et al. (2023).
@@
UQ methods can also be classified into white-box and black-box approaches depending on the required access to LLM itself and its outputs. Black-box techniques do not require any other input except the model outputs.
@@
The method in our work can be attributed to the information-based group, and can be applied only to white-box LLMs because it requires access to the probability distribution of the generated tokens. Compared to other techniques, it offers a novel approach to post-processing the probability distribution and is designed to quantify uncertainty in parts of the output, such as atomic claims and individual words.
@@ 3 Fact-Checking Pipeline
The fact-checking pipeline (see Figure 5) starts with splitting a generated text into atomic claims, e.g. using a much smaller model fine-tuned for this particular task. For experimental evaluations in this work, we follow the approach of FactScore Min et al. (2023), where splitting is implemented via the ChatGPT API.
Each atomic claim is matched against the sequence of tokens in the original text with corresponding probability distributions. Then we calculate token-level uncertainty scores and aggregate them into the claim-level uncertainty.
@@
Figure 2: Example of CCP calculation on the Vicuna 13b generation for the word painting.
@@ 4 Uncertainty Quantification
4.1 Claim-Level Uncertainty Quantification Baselines
@@
$MP(C)=1-\prod_{j\in C}P(x_{j}\mid x_{<j}).$ (1)
@@
P(True), similar to Kadavath et al. (2022), measures the uncertainty of the claim by asking the LLM itself whether the generated claim is true or not, as calculated by the model confidence of the first generated token $y_{1}$ being equal to “True”:
@@ 4.2.1 Motivation and Theoretical Background
When a LLM generates an output, it faces various types of uncertainty reflected in the token distribution of the current generation step (see Figure 4 for an example). We identify three distinct types of uncertainty:
  1. 1.
    Claim type/order uncertainty: What claim to generate on the current step? For example, on the current step, a LLM might hesitate between generating a year of graduation of a person and a field of study. A different order of claims, missing claims, or different types of generated claims do not make produced text less factual. Therefore, when performing fact-checking, we should not take this type of uncertainty into account.
  2. 2.
    Surface form uncertainty: What synonyms or hypernyms to use when generating a claim (e.g. “art” or “painting”)? Different surface forms also do not make the text less factual — they might change the style, but not the underlying meaning of the text. Therefore, this type of uncertainty is also not relevant for fact-checking.
  3. 3.
    Claim uncertainty: What specific piece of information to relay for a particular claim type? For example, an LLM might not be sure what field of study to generate, producing a token distribution with multiple highly-probable variants, such as “painting”, “acting”, “sculpture”. In a similar way, for the year of graduation, a LLM might produce a distribution with various potential years. This uncertainty is relevant for fact-checking, because if the model is not sure about the information it relays, there might be a high chance of a factual mistake.
@@
$CCP(x_{j})=P\bigl(\text{Meaning}(x_{1:j})\mid x_{<j},\text{ClaimType}(x_{1:j})\bigr).$ (4)
@@
The meaning function and the construction of the claim type set of words can be implemented in various ways. We outline one approach in Section˜4.2.2.
@@ 4.2.2 Implementation
The distribution $X_{j}$ at the position $j$ is approximated by top-$K$ alternatives $\left\{x^{k}_{j}\right\}_{k=1}^{K}$ with $x^{1}_{j}\equiv x_{j}$. We replace $x_{j}$ with its alternatives $x_{j}^{k}$ and obtain new instances $x_{<j}\circ x^{k}_{j},k=1,\dots,K$. Each new instance is compared against the original prediction $x_{1:j}=x_{<j}\circ x_{j}$ using an NLI model. We define $\texttt{NLI}(x_{j}^{k},x_{j}):=\texttt{NLI}(x_{<j}\circ x^{k}_{j},x_{1:j})$, where $\texttt{NLI}(x_{<j}\circ x^{k}_{j},x_{1:j})$ means application of the NLI model to the sentences $x_{<j}\circ x^{k}_{j}$ and $x_{1:j}$. The outcome of NLI procedure is one of three classes: entail (‘e’), contradict (‘c’) or neutral (‘n’). If the new instance entails the original prediction $\texttt{NLI}(x_{j}^{k},x_{j})=\textrm{`e'}$, then we consider $x_{<j}\circ x^{k}_{j}$ has the same meaning with $x_{1:j}$ ($x^{k}_{j}\in M(x_{j})$) and corresponds to the same claim type ($x^{k}_{j}\in CT(x_{j})$). If the new instance contradicts the original prediction $\texttt{NLI}(x_{j}^{k},x_{j})=\textrm{`c'}$, then we consider $x_{<j}\circ x^{k}_{j}$ has a different meaning with $x_{1:j}$ ($x^{k}_{j}\notin M(x_{j})$), but corresponds to the same claim type ($x^{k}_{j}\in CT(x_{j})$). Otherwise, if the new instance is neutral w.r.t. the original prediction $\texttt{NLI}(x_{j}^{k},x_{j})=\textrm{`n'}$, then we consider that $x_{<j}\circ x^{k}_{j}$ does not correspond to the same claim type as $x_{1:j}$ ($x^{k}_{j}\notin CT(x_{j})$). Thus, Equation (5) for CCP can be written as follows:
$CCP_{\textit{word}}(x_{j})=\dfrac{\sum_{k:\texttt{NLI}(x^{k}_{j},x_{j})=\texttt{`e'}}P(x^{k}_{j}\mid x_{<j})}{\sum_{k:\texttt{NLI}(x^{k}_{j},x_{j})\in\left\{\texttt{`e'},\texttt{`c'}\right\}}P(x^{k}_{j}\mid x_{<j})}.$ (6)
@@
We note that most transformer LLMs generate sub-word tokens instead of whole words. To obtain the distribution for whole words, we generate one or multiple tokens using beam search with $K$ beams.
@@
An example of calculating the CCP for a fact is presented in Figure 2. Other detailed examples of CCP calculation are available in Appendix A.

### v2

@@ Abstract
Large language models (LLMs) are notorious for hallucinating, i.e., producing erroneous claims in their output. Such hallucinations can be dangerous, as occasional factual inaccuracies in the generated text might be obscured by the rest of the output being generally factually correct, making it extremely hard for the users to spot them. Current services that leverage LLMs usually do not provide any means for detecting unreliable generations. Here, we aim to bridge this gap. In particular, we propose a novel fact-checking and hallucination detection pipeline based on token-level uncertainty quantification. Uncertainty scores leverage information encapsulated in the output of a neural network or its layers to detect unreliable predictions, and we show that they can be used to fact-check the atomic claims in the LLM output. Moreover, we present a novel token-level uncertainty quantification method that removes the impact of uncertainty about what claim to generate on the current step and what surface form to use. Our method Claim Conditioned Probability (CCP) measures only the uncertainty of a particular claim value expressed by the model. Experiments on the task of biography generation demonstrate strong improvements for CCP compared to the baselines for seven LLMs and four languages. Human evaluation reveals that the fact-checking pipeline based on uncertainty quantification is competitive with a fact-checking tool that leverages external knowledge.
@@ 1 Introduction
Large language models (LLMs) have become a ubiquitous and versatile tool for addressing a variety of natural language processing (NLP) tasks. People use these models for tasks including information search Sun et al. (2023b), to ask medical questions Thirunavukarasu et al. (2023), or to generate new content Sun et al. (2023a). Recently, there has been a notable shift in user behavior, indicating an increasing reliance on and trust in LLMs as primary information sources, often surpassing traditional channels. However, a significant challenge with the spread of these models is their tendency to produce “hallucinations”, i.e., factually incorrect generations that contain misleading information Bang et al. (2023); Dale et al. (2023). This is a side-effect of the way modern LLMs are designed and trained Kalai and Vempala (2023).
LLM hallucinations are a major concern because the deceptive content at the surface level can be highly coherent and persuasive. Common examples include the creation of fictitious biographies or the assertion of unfounded claims. The danger is that a few occasional false claims might be easily obscured by a large number of factual statements, making it extremely hard for people to spot them. As hallucinations in LLM outputs are hard to eliminate completely, users of such systems could be informed via highlighting some potential caveats in the text, and this is where our approach can help.
Fact-checking is a research direction that addresses this problem. It is usually approached using complex systems that leverage external knowledge sources Guo et al. (2022); Nakov et al. (2021); Wadden et al. (2020). This introduces problems related to the incomplete nature of such sources and notable overhead in terms of storing the knowledge. We argue that information about whether a generation is a hallucination is encapsulated in the model output itself, and can be extracted using uncertainty quantification (UQ) Gal et al. (2016); Kotelevskii et al. (2022); Vazhentsev et al. (2022, 2023a). This avoids implementing complex and expensive fact-checking systems that require additional computational overhead and rely on external resources.
Prior work has mainly focused on quantification of uncertainty for the whole generated text and been mostly limited to tasks such as machine translation Malinin and Gales (2020), question answering Kuhn et al. (2023), and text summarization van der Poel et al. (2022). However, the need for an uncertainty score for only a part of the generation substantially complicates the problem. We approach it by leveraging token-level uncertainty scores and aggregating them into claim-level scores. Moreover, we introduce a new token-level uncertainty score, namely claim-conditioned probability (CCP), which demonstrates confident improvements over several baselines for seven LLMs and four languages.
@@
  • We propose a novel method for token-level uncertainty quantification that outperforms baselines and can be used as a plug-in in a fact-checking framework.
@@
  • We provide an empirical and ablation analysis of the method for fact-checking of LLM generations, and find that the uncertainty scores we produce can help to spot claims with factual errors for seven LLMs over four languages: English, Chinese, Arabic, and Russian.
  • The method is implemented as a part of the LM-Polygraph library Fadeeva et al. (2023). All the code and data for experiments is publicly available [^1] https://github.com/IINemo/lm-polygraph.
@@
Figure 1: Visual comparison of our Claim-Conditioned Probability method to the Maximum Probability baseline. CCP accurately identifies the incorrectly specified number of awards (in red), whereas Maximum Probability erroneously highlights the claim that is actually correct.
@@ 2.1 Fact-Checking LLM Generations and Detecting Hallucinations
The problem of hallucinations has made fact-checking of LLM outputs a prominent topic in the research community and resulted in a surge of publications on the topic. Chern et al. (2023) present Factool – a task and domain agnostic framework for hallucination detection that leverages GPT for claim extraction and verification. Manakul et al. (2023) suggest to sample multiple outputs from black-box LLMs and evaluate how similar the sampled responses are using the external model. Varshney et al. (2023) detect LLM hallucinations by extracting key parts of output using an external model and estimating their uncertainty based on logits. The most uncertain parts are verified using an external knowledge source. Pan et al. (2023) propose to fact-check complex statements by decomposing them into simpler subtasks and generating reasoning programs to verify these statements. Min et al. (2023) present a methodology for evaluating long LLM-generated texts by decomposing them into simple atomic statements and further verifying them against some knowledge source. Several subsequent works further optimize components of the knowledge-based fact-checking pipelines (Wang et al., 2024a, b).
@@
In contrast to previous work that leverages external knowledge sources for fact-checking (a database or another LLM), our work is the first to investigate token-level UQ methods for this task using the LLM and its outputs only.
2.2 Uncertainty Quantification of LLM Generations
@@
UQ techniques for LLM generation can be classified into five major categories Fadeeva et al. (2023). Information-based methods leverage the probability distribution of generated tokens and usually do not require any additional models. In this category, we can include methods such as perplexity Fomicheva et al. (2020), mean token entropy Fomicheva et al. (2020), point-wise mutual information (PMI) Takayama and Arase (2019), and conditional PMI van der Poel et al. (2022).
@@
Another category of methods is based on density estimation of latent instance representations. A typical example in this category is Mahalanobis distance Lee et al. (2018) and its various modifications Ren et al. (2023); Vazhentsev et al. (2023b). The disadvantage of such methods is the need for access to the LLM training data in order to fit external density models, which is problematic for most general-purpose LLMs.
@@
Ensembling and Monte Carlo dropout methods are based on the lexical diversity of multiple outputs sampled from one or multiple versions of LLMs for a single query Malinin and Gales (2020); Fomicheva et al. (2020). Their main drawback is that they require many predictions, which makes them too computationally and memory intensive for practical purposes. Additionally, it is difficult to apply these techniques to quantify uncertainty of text fragments such as claims, as different samples may seriously diverge.
@@
It has recently been shown that LLMs can reflexively estimate the confidence of their generations simply by asking themselves about the truthfulness of their output (Kadavath et al., 2022). This can work better than analyzing the probability distribution for the original prediction, but requires a second pass of inference, feeding the original output as a part of the query.
@@
Finally, there is a group of methods that leverages the diversity of meanings that the LLM generates for a given query. This group includes semantic entropy Kuhn et al. (2023) and various scores based on the analysis of the similarity matrix between outputs Lin et al. (2023).
@@
UQ methods can also be classified into white-box and black-box Lin et al. (2023) approaches, depending on the required access to LLM itself and its outputs. Black-box techniques do not require any other input except generated texts.
@@
The method in our work can be attributed to the information-based group, and can be applied only to white-box LLMs because it requires access to the probability distribution of the generated tokens. Compared to other techniques, it offers a novel approach to post-processing the probability distribution and is specifically designed to quantify uncertainty of output fragments, such as atomic claims and individual words.
@@ 3 Fact-Checking Pipeline
The fact-checking pipeline (see Figure 4) starts with splitting a generated text into atomic claims, e.g. using a much smaller model fine-tuned for this particular task. For experimental evaluations in this work, we follow the FactScore approach Min et al. (2023), where splitting is implemented via the OpenAI Chat API.
Each atomic claim is matched against the sequence of tokens in the original text with the corresponding probability distributions. Then we calculate token-level uncertainty scores and aggregate them into the claim-level uncertainty.
@@
Figure 2: Example of CCP calculation for the word painting in a Vicuna 13b generation.
@@ 4 Uncertainty Quantification
4.1 Claim-Level UQ Baselines
@@
$MP(C)=1-\prod\nolimits_{j\in C}P(x_{j}\mid x_{<j}).$ (1)
Perplexity is a common metric used to evaluate the performance of LLMs. Lower perplexity indicates that a model’s probability distribution better predicts a sample. It is computed as the average negative log probability of generated tokens that belong to the claim $C$:
$Perp(C)=\exp\left(-\frac{1}{|C|}\sum\nolimits_{j\in C}\log P(x_{j}\mid x_{<j})\right).$ (2)
@@
P(True), similar to Kadavath et al. (2022), measures the uncertainty of the claim by asking the LLM itself whether the generated claim is true or not. The confidence is the probability of the first generated token $y_{1}$ being equal to “True”:
@@ 4.2.1 Motivation and Theoretical Background
When an LLM generates an output, it faces various types of uncertainty reflected in the token distribution of the current generation step (see Figure 5 for an example). We identify three distinct types of uncertainty:
(1) Claim type/order uncertainty: What claim to generate on the current step? For example, on the current step, an LLM might hesitate between generating a year of graduation of a person and a field of study. A different order of claims, missing claims, or different types of generated claims do not make produced text less factual. Therefore, when performing fact-checking, we should not take this type of uncertainty into account.
(2) Surface form uncertainty: What synonyms or hypernyms to use when generating a claim (e.g. “art” or “painting”)? Different surface forms also do not make the text less factual — they might change the style, but not the underlying meaning of the text. Therefore, this type of uncertainty is also not relevant for fact-checking.
(3) Claim uncertainty: What specific piece of information to relay for a particular claim type? For example, an LLM might not be sure which field of study to generate, producing a token distribution with multiple highly-probable variants, such as “painting”, “acting”, “sculpture”. Similarly, for the year of graduation, an LLM might produce a distribution with various potential years. This uncertainty is relevant for fact-checking, because if the model is not sure about the information it relays, there might be a high chance of a factual mistake.
@@
$CCP(x_{j})=P\bigl(\text{Meaning}(x_{1:j})|x_{<j},\text{ClaimType}(x_{1:j})\bigr).\!\!\!$ (5)
@@
The meaning function and the construction of the set of words that belong to the same claim type can be implemented in various ways. We outline our approach in Section˜4.2.2.
@@ 4.2.2 Implementation
The distribution $X_{j}$ at the position $j$ is approximated by top-$K$ alternatives $\{x^{k}_{j}\}_{k=1}^{K}$ with $x^{1}_{j}\equiv x_{j}$. We replace $x_{j}$ with its alternatives $x_{j}^{k}$ and obtain new instances $x_{<j}\circ x^{k}_{j},k=1,\dots,K$. Each new instance is compared against the original prediction $x_{1:j}=x_{<j}\circ x_{j}$ using an NLI model. We define $\texttt{NLI}(x_{j}^{k},x_{j}):=\texttt{NLI}(x_{<j}\circ x^{k}_{j},x_{1:j})$, where $\texttt{NLI}(x_{<j}\circ x^{k}_{j},x_{1:j})$ means application of the NLI model to the text fragments $x_{<j}\circ x^{k}_{j}$ and $x_{1:j}$.
The outcome of the NLI procedure is one of three labels: entail (‘e’), contradict (‘c’), or neutral (‘n’). If the new instance entails the original prediction $\texttt{NLI}(x_{j}^{k},x_{j})=\textrm{`e'}$, then we consider $x_{<j}\circ x^{k}_{j}$ has the same meaning with $x_{1:j}$ ($x^{k}_{j}\in M(x_{j})$) and corresponds to the same claim type ($x^{k}_{j}\in CT(x_{j})$).
If the new instance contradicts the original prediction $\texttt{NLI}(x_{j}^{k},x_{j})=\textrm{`c'}$, then we consider $x_{<j}\circ x^{k}_{j}$ has a different meaning with $x_{1:j}$ ($x^{k}_{j}\notin M(x_{j})$), but corresponds to the same claim type ($x^{k}_{j}\in CT(x_{j})$). Otherwise, if the new instance is neutral w.r.t. the original prediction $\texttt{NLI}(x_{j}^{k},x_{j})=\textrm{`n'}$, then we consider that $x_{<j}\circ x^{k}_{j}$ does not correspond to the same claim type as $x_{1:j}$ ($x^{k}_{j}\notin CT(x_{j})$). Thus, equation (6) for CCP can be written as follows:
$CCP_{\textit{word}}(x_{j})=\dfrac{\sum_{k:\texttt{NLI}(x^{k}_{j},x_{j})=\texttt{`e'}}P(x^{k}_{j}\mid x_{<j})}{\sum_{k:\texttt{NLI}(x^{k}_{j},x_{j})\in\left\{\texttt{`e'},\texttt{`c'}\right\}}P(x^{k}_{j}\mid x_{<j})}.$
@@
We note that most transformer LLMs generate sub-word tokens instead of whole words. To obtain distributions for whole words, we generate one or multiple tokens using beam search with $K$ beams.
@@
An example of calculating the CCP for a claim is presented in Figure 2. Other detailed examples of CCP calculation are available in Appendix B.